# From Medical Records to Diagnostic Dialogues: A Clinical-Grounded Approach and Dataset for Psychiatric Comorbidity

### v1

@@ 1 Introduction
Psychiatric disorders account for over 125 million disability-adjusted life years globally (Collaborators et al., 2022). A major challenge lies in psychiatric comorbidity, i.e., the co-occurrence of multiple conditions, which significantly complicates diagnosis and treatment. For instance, in a Netherlands study of depression and anxiety, 67% of individuals with a primary depression diagnosis had current and 75% had lifetime comorbid anxiety disorders (Lamers et al., 2011). Yet, most existing datasets and models narrowly focus on single disorders (Aich et al., 2024; Yao et al., 2022), while the few corpora covering multiple conditions lack fine-grained annotations and fail to capture symptom co-occurrence and progression within diagnostic processes (Sun et al., 2021; Yin et al., 2025; Cohan et al., 2018). As a result, large language models (LLMs) have not been systematically evaluated on multi-disorder diagnostic tasks, limiting the development of reliable screening systems that require step-by-step reasoning grounded in DSM-5 standards (Association et al., 2013).  [^1] DSM-5 is the fifth edition of the Diagnostic and Statistical Manual of Mental Disorders by the American Psychiatric Association. It provides standardized criteria for diagnosing mental disorders and is used to ensure accuracy and consistency.
@@ 3.1 EMR Generation
Generation Strategy  We adopted a modular approach to EMR generation, which outperformed a single-step aggregation method in terms of information recall, classification accuracy, and reasoning coherence. This approach involves categorizing post content into the seven standard EMR sections, generating each section individually using either rule-based methods or LLM inference, and merging them into a complete record.
@@ 3.2 PsyCoProfile Dataset description and statistics
To validate the realism of the synthetic EMRs, we compared them with real-world data [^2] For real-world comparison, we obtained approximately 1,000 de-identified clinical records from the National Mental Health Center, provided with institutional approval and used strictly for research purposes. across disease distribution, demographic factors (age and gender), and family history. The results are illustrated in Figure 3. Disease prevalence in synthetic EMRs broadly aligns with real-world data while exhibiting a more balanced distribution. Age shows a peak in the 20–24 group for synthetic data versus 30–34 for real records (Collaborators et al., 2022), likely reflecting the younger social media demographic (DataReportal, 2020). Gender proportions are more balanced in synthetic records, consistent with reduced stigma in online contexts (Porteous & Armstrong, 2021). Family history is slightly overrepresented but remains within clinically plausible bounds. Psychiatrist feedback on sampled records confirmed sufficient diversity and clinical plausibility, closely matching real-world documentation.
@@ 4.1 Clinical-Grounded Interview Pipeline
The Structured Clinical Interview for DSM-5 (SCID-5-RV) is a standardized tool for assessing mental disorders, offering a systematic approach to symptom evaluation and diagnosis. Adhering to SCID-5-RV grounds our framework in established clinical practices. To guide LLM-based psychiatric interviews, we restructure SCID-5-RV into two parts: a Hierarchical Diagnostic State Machine (HDSM) and a Diagnostic Context Tree (DCT). This design ensures a structured, dynamic, and clinically coherent dialogue flow.
@@
Binary symptom scale and flow control. The original four-point SCID-5-RV symptom scale (“unclear”, “absent”, “subthreshold”, “threshold”) is reduced to a binary form: present or absent. Colored arrows in Figure 5 illustrate how different responses guide the interview trajectory. If a BLS node has no valid successor, the agent performs a localized random traversal within the sub-state group. The transition to the next diagnostic state occurs only after all terminal BLS nodes (i.e., those without successors) have been visited. If the number of “positive” (i.e., “present”) responses exceeds a threshold, the group is deemed positive and follows the corresponding path; otherwise, it is absent and triggers an alternative transition.
@@ 5 Final Product: PsyCoTalk Dataset
Evaluation  We conduct two types of human evaluation. First, 50 dialogues from PsyCoTalk are rated in a double-blind fashion by five licensed psychiatrists with more than seven years of clinical experience. Each dialogue is evaluated across six dimensions grouped under four criteria: Professionalism, Communication, Fluency, and Realism. Professionalism assesses whether the doctor successfully elicits all symptoms necessary for diagnosis. Communication evaluates (i) the doctor’s proactivity in questioning and (ii) the patient’s responsiveness. Fluency measures (i) syntactic and topical coherence and (ii) avoidance of redundancy. Realism assesses how closely the dialogue resembles actual psychiatric consultations. Second, we perform an AB test using 10 two-turn excerpts sampled from each dataset (PsyCoTalk, D4, MDD-5k, and Real-World Dial). The same five psychiatrists are asked to judge whether each excerpt appears real or AI-generated. Each excerpt identified as “real” receives one point, and final realism scores are normalized to a 10-point scale.
@@
To assess diagnostic accuracy, we use initial EMR labels as ground truth and evaluate with exact-match (all 5 disorders correct). A zero-shot Qwen2.5-72B baseline reaches 0.22 subset accuracy, while our HDSM-guided multi-agent system improves to 0.31, a significant gain (McNemar’s test, $p=7\times 10^{-6}<0.001$). On 200 sampled cases, GPT-4o-mini and Deepseek-v3 achieve below 0.1, Qwen3-32B below 0.02, and Qwen3-8B below 0.04. Per-label F1: MDD (0.92), AD (0.81), ADHD (0.64), BD (0.40). These align with clinical trends where BD and ADHD are harder to diagnose than MDD and AD due to symptom overlap, subtler onset, and higher heterogeneity (Hui et al., 2018; Barkley & Brown, 2008).
@@
PsyCoTalk focuses on four prevalent psychiatric disorders and their comorbidities, which reflects common real-world cases but limits coverage of rarer conditions. While we conducted a small-scale English generation experiment, the main dataset remains Chinese, restricting multilingual applicability. Nevertheless, our pipeline is extensible and can be scaled to broader disorder coverage and cross-lingual settings in future work.
@@ 6 Conclusion
The authors have read and adhere to the ICLR Code of Ethics. This work does not involve human subjects, identifiable private data, or harmful applications. No external sponsorship or conflict of interest influenced the design or conclusions of this work.
@@
Intended use and risk mitigation. The dataset is designed for research on multi-disorder screening and dialogue modelling. It is not a diagnostic tool, nor does it provide treatment advice. Synthetic conversations may still reflect biases inherited from web data and large language models; users must apply rigorous evaluation before any downstream use. We urge practitioners to keep humans in the loop when analysing sensitive mental-health text, and to follow relevant professional, legal and ethical guidelines. Future work will extend coverage while continuing to consult ethics boards and domain experts.
@@ A.2.2 Comparison with Real Clinical Dialogues
Metric
PsyCoTalk
(Chinese, 3k)
Real Clinical
(Chinese, 1.7k)
PsyCoTalk-eng
AlexanderStreet
(English, 1.3k)
@@ A.2.4 Word Frequency and Part-of-Speech Comparison
Word Class
Chinese
High-freq
Count (%)
English
High-freq
Count (%)
Chinese
Total Freq (%)
English
Total Freq (%)

### v2

@@ 1 Introduction
Psychiatric disorders account for over 125 million disability-adjusted life years globally (Collaborators and others, 2022). A major challenge lies in psychiatric comorbidity, i.e., the co-occurrence of multiple conditions, which significantly complicates diagnosis and treatment. For instance, in a Netherlands study of depression and anxiety, 67% of individuals with a primary depression diagnosis had current and 75% had lifetime comorbid anxiety disorders (Lamers et al., 2011). Yet, most existing datasets and models narrowly focus on single disorders (Aich et al., 2024; Yao et al., 2022), while the few corpora covering multiple conditions lack fine-grained annotations and fail to capture symptom co-occurrence and progression within diagnostic processes (Sun et al., 2021; Yin et al., 2025; Cohan et al., 2018). As a result, large language models (LLMs) have not been systematically evaluated on multi-disorder diagnostic tasks, limiting the development of reliable screening systems that require step-by-step reasoning grounded in DSM-5 standards (Association and others, 2013).  [^1] DSM-5 is the fifth edition of the Diagnostic and Statistical Manual of Mental Disorders by the American Psychiatric Association. It provides standardized criteria for diagnosing mental disorders and is used to ensure accuracy and consistency.
@@ 3.1 EMR Generation
Generation Strategy  We adopted a modular approach to EMR generation, which outperformed a single-step aggregation method in terms of information recall, classification accuracy, and reasoning coherence. This approach involves categorizing post content into the seven standard EMR sections, generating each section individually using either rule-based methods or LLM inference, and merging them into a complete record. During EMR generation, we iteratively sampled cases for expert review by eight psychiatrists; their feedback guided revisions to ensure medical validity.
@@ 3.2 PsyCoProfile Dataset description and statistics
To validate the realism of the synthetic EMRs, we compared them with real-world data [^2] For real-world comparison, we obtained approximately 1,000 de-identified clinical records from the National Mental Health Center, provided with institutional approval and used strictly for research purposes. across disease distribution, demographic factors (age and gender), and family history. The results are illustrated in Figure 3. Disease prevalence in synthetic EMRs broadly aligns with real-world data while exhibiting a more balanced distribution. Age shows a peak in the 20–24 group for synthetic data versus 30–34 for real records (Collaborators and others, 2022), likely reflecting the younger social media demographic (DataReportal, 2020). Gender proportions are more balanced in synthetic records, consistent with reduced stigma in online contexts (Porteous and Armstrong, 2021). Family history is slightly overrepresented but remains within clinically plausible bounds. Psychiatrist feedback on sampled records confirmed sufficient diversity and clinical plausibility, closely matching real-world documentation.
@@ 4.1 Clinical-Grounded Interview Pipeline
The Structured Clinical Interview for DSM-5 (SCID-5-RV) is a standardized tool for assessing mental disorders, offering a systematic approach to symptom evaluation and diagnosis. Its diagnostic structure has been widely validated in diverse clinical and cultural settings (Mohammadkhani et al., 2020; Shankman et al., 2018; Brodey et al., 2018). Adhering to SCID-5-RV grounds our framework in established clinical practices. To guide LLM-based psychiatric interviews, we restructure SCID-5-RV into two parts: a Hierarchical Diagnostic State Machine (HDSM) and a Diagnostic Context Tree (DCT). This design ensures a structured, dynamic, and clinically coherent dialogue flow.
@@
Binary symptom scale and flow control. The original four-point SCID-5-RV symptom scale (“unclear”, “absent”, “subthreshold”, “threshold”) is reduced to a binary form: present or absent. We do not discard severity information in this simplification; it is encoded in specific HDSM nodes and transitions (e.g., duration constraints such as A254 $\geq$ 1 week, and functional impairment checks such as A23 on school, work, or social functioning), while the binary variables only indicate whether diagnostic thresholds are met. We adopt this binarization not because the four-level severity scale is unimportant, but because distinctions such as “below threshold” versus “insufficient information” require nuanced clinical judgment: in our early experiments, asking LLM agents to maintain four-level severity ratings frequently produced inconsistent or contradictory outputs, leading to diagnostic loops or invalid transitions, whereas binary thresholding yielded a more stable and auditable diagnostic process. Colored arrows in Figure 5 illustrate how different responses guide the interview trajectory. If a BLS node has no valid successor, the agent performs a localized random traversal within the sub-state group. The transition to the next diagnostic state occurs only after all terminal BLS nodes (i.e., those without successors) have been visited. If the number of “positive” (i.e., “present”) responses exceeds a threshold, the group is deemed positive and follows the corresponding path; otherwise, it is absent and triggers an alternative transition.
@@ 5 Final Product: PsyCoTalk Dataset
[…]
@@
To assess diagnostic accuracy, we use initial EMR labels as ground truth and evaluate with exact-match (all 5 disorders correct). A zero-shot Qwen2.5-72B baseline reaches 0.22 subset accuracy, while our HDSM-guided multi-agent system improves to 0.31, a significant gain (McNemar’s test, $p=7\times 10^{-6}<0.001$). On 200 sampled cases, GPT-4o-mini and Deepseek-v3 achieve below 0.1, Qwen3-32B below 0.02, and Qwen3-8B below 0.04. Per-label F1: MDD (0.92), AD (0.81), ADHD (0.64), BD (0.40). These align with clinical trends where BD and ADHD are harder to diagnose than MDD and AD due to symptom overlap, subtler onset, and higher heterogeneity (Hui et al., 2018; Barkley and Brown, 2008).
@@
PsyCoTalk focuses on four prevalent psychiatric disorders and their comorbidities, which reflects common real-world cases but limits coverage of rarer conditions. This choice is driven by the scarcity of clinically reliable comorbidity data, so we retain only combinations with sufficient symptom diversity and sample size to support stable EMR construction. While we conducted a small-scale English generation experiment, the main dataset remains Chinese due to expert-evaluation constraints, restricting multilingual applicability. Moreover, Reddit-derived symptom cues may introduce demographic or cultural skew, although the EMR template and SCID-5-based HDSM are language- and culture-agnostic and can be ported to local clinical corpora. Nevertheless, our pipeline is extensible and can be scaled to broader disorder coverage and cross-lingual settings. Future work will conduct more experiments and evaluations to quantify the dataset’s impact on model performance and diagnostic reasoning.
@@ 6 Conclusion
The authors have read and adhere to the ICLR Code of Ethics. This work does not involve human subjects, identifiable private data, or harmful applications. No real patient data are used; the only human involvement is expert evaluation by five licensed psychiatrists, who provided informed consent and were compensated according to responsible-research guidelines. No external sponsorship or conflict of interest influenced the design or conclusions of this work.
@@
Intended use and risk mitigation. The dataset is designed for research on multi-disorder screening and dialogue modelling. It is not a diagnostic tool, nor does it provide treatment advice or actionable clinical guidance. To prevent misuse, we define compliant use boundaries as follows: (1) Research-only restriction: all EMRs and dialogues are synthetic and must not be used for individual clinical decisions, real-world psychological counseling, or any form of patient-facing service; (2) Data-usage agreement: access requires signing an agreement that prohibits clinical deployment, real-user psychological services, and any re-identification attempts. Any real-world mental-health use of models trained on this dataset would require formal ethics review and clinician oversight. Synthetic conversations may still reflect biases inherited from web data and large language models; users must apply rigorous evaluation before any downstream use. We urge practitioners to keep humans in the loop when analysing sensitive mental-health text, and to follow relevant professional, legal and ethical guidelines. Future work will extend coverage while continuing to consult ethics boards and domain experts.
@@
Acknowledgement
This work has been supported by the China NSFC Project (No. 62572320), the Shanghai Shenkang Hospital Development Center Project (Grant No. SHDC12025118), Tianqiao and Chrissy Chen Institute (TCCI) and EverMind AI, Inc (Program: Mental Health Data Synthesis and Scientific Data Competition).
Reproducibility Statement
We have made all code and datasets used in this work publicly available in an anonymous repository to ensure full reproducibility of our experiments: https://github.com/X-LANCE/PsyCo-Diagnosis.
@@ A.2.2 Comparison with Real Clinical Dialogues
Metric PsyCoTalk (Chinese, 3k) Real Clinical (Chinese, 1.7k) PsyCoTalk-eng (205) AlexanderStreet (English, 1.3k)
@@ A.2.4 Word Frequency and Part-of-Speech Comparison
Word Class Chinese High-freq Count (%) English High-freq Count (%) Chinese Total Freq (%) English Total Freq (%)